Determine the actual Qerror(mV) of PCM with 400mV resolution for sample voltage of 2.3V use 12
bit
ANS: 100
Answer #1

Solution for the problem is provided below:

It is given that;

            The resolution of PCM, dv = 400 mV

            Sample voltage = 2.3 V

            Number of bits used = 12

The length of the quantified sample = 212 = 4096

To find the output voltage range, mp = (dv * L)/2 = (400 * 4096/2) = 819200

Actual Qerror = (mp/2L )= 819200/(2*4096) = 100

Thus actual Qerror = 100